OpenStudy (anonymous):

What is the next number in the sequence 1, 8, 27, 64

OpenStudy (anonymous):

1=1^3 8=2^3 27=3^3 64=4^3

OpenStudy (anonymous):

So what do u think will be the next number

OpenStudy (anonymous):

@CaptainKay ???